Assume that the cost of investment is 10 000 and will result in cash flows of 6 000 in the first year and 6 000 in year 2.the companys cost of capital is 20%.find the net present value.
Answer by Theo(13342)

you have -10000 in time point 0
you have +6000 in time point 1
you have +6000 in time point 2

cost of capital is 20%.

discount the cash flow to time point 0 at 20%.

-1000 / 1.20 ^ 0 = -10000
+6000 / 1.20 ^ 1 = +5000
+6000 / 1.20 ^ 2 = 4166.67

add up the present values to get -833.33.

that's your net present value.